# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ V-D With Large Language Models
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [148], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning [149, 150] is a key step in aligning LLM with more desired behaviors, but requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [151] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Some other work [152, 153] also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. SPAG [154] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on a wide range of reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero [155] which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work [156] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research efforts.
V-E Reality Applications
@@
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities [161], and to address the image retargeting problem [162]. Moreover, EvoPlay [163] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [164] to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for adversarial catching tasks, with significant efforts dedicated to Sim2Real transitions for real-world success [165].
VI Conclusion
The burgeoning field of self-play in RL has been systematically explored in this survey. The core idea of self-play, in which agents interact with copies or past versions of themselves, has proven to be a powerful approach for developing robust strategies across various domains. Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the background of self-play. Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Details are meticulously provided on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities. The transition from theory to application is underscored by a rigorous analysis of self-play’s role within complex scenarios, such as board games, card games, and video games. Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands. Future work may focus on solving these problems, integrating self-play with LLMs, and achieving real-world applications. In conclusion, self-play stands as a cornerstone of modern RL research, offering profound insights and tools for developing advanced AI systems. This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this dynamic and evolving field.
@@ Acknowledgment
This research was supported by National Natural Science Foundation of China (No.62325405, U19B2019, M-0248), Tsinghua University Initiative Scientific Research Program, Tsinghua-Meituan Joint Institute for Digital Life, Beijing National Research Center for Information Science, Technology (BNRist) and Beijing Innovation Center for Future Chips, Postdoctoral Fellowship Program of CPSF under Grant Number GZC20240830, China Postdoctoral Science Special Foundation 2024T170496.

### v2

@@ 5.4 With Large Language Models
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [1], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning [102, 3] is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [26] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Another line of work [98, 121, 135] formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game. Some other work [56, 146] also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. Self-Play of Adversarial Game (SPAG) [27] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero [38] which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work [138] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
5.5 Real-World Applications
@@
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities [59], and to address the image retargeting problem [66]. Moreover, EvoPlay [132] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [65] to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football [81, 47], quadruped competition [136], and multi-drone volleyball [139], with significant efforts dedicated to Sim2Real transitions for real-world success [41].
6 Conclusion
The burgeoning field of self-play in RL has been systematically explored in this survey. The core idea of self-play that agents interact with copies or past versions of themselves has proven to be a powerful approach for developing robust strategies across various domains. Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the game theory background of self-play. Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. We provide detailed explanations on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities. The transition from theory to application is underscored by a rigorous analysis of self-play’s role within various scenarios, such as board games, card games, and video games. Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands. Also, future work can focus on how to integrate self-play with LLMs and achieve real-world applications. In conclusion, self-play is vital to modern RL research, offering key insights and tools for developing advanced AI systems. This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this field.
@@ Acknowledgment
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
Appendix A Appendix
A.1 Abbreviations in Alphabetical Order
ABR Approximate Best Response
AI Artificial Intelligence
ARM Advantage-based Regret Minimization
ARMAC Advantage Regret-Matching Actor-Critic
BR Best Response
CCE Coarse Correlated Equilibrium
CE Correlated Equilibrium
CFR Counterfactual Regret Minimization
COP Combinatorial Optimization Problem
CSPL Curriculum Self-Play Learning
CVRP Capacitated Vehicle Routing Problem
DCFR Discounted Counterfactual Regret Minimization
DO Double Oracle
EGTA Empirical Game-Theoretic Analysis
FP Fictitious Play
FSP Fictitious Self-Play
FTW For The Win
FXP Fictitious Cross-Play
GAIL Generative Adversarial Imitation Learning
GRF Google Research Football
HULHE Heads-Up Limit Texas Hold’em
HUNL Heads-Up No-Limit Texas Hold’em
JPSRO Joint Policy-Space Response Oracle
LCFR Linear Counterfactual Regret Minimization
LLM Large Language Model
MARL Multi-Agent Reinforcement Learning
MCCFR Monte Carlo Counterfactual Regret Minimization
MCTS Monte Carlo Tree Search
MDP Markov Decision Process
MG Markov Game
ML Machine Learning
MOBA Multiplayer Online Battle Arena
MSS Meta-Strategy Solver
NE Nash Equilibrium
NeuPL Neural Population Learning
NFSP Neural Fictitious Self-Play
ODO Online Double Oracle
PBR Preference-based Best Response
PBS Public Belief State
PFSP Prioritized Fictitious Self-Play
POMG Partially Observable Markov Game
PSRO Policy-Space Response Oracle
RCFR Regression Counterfactual Regret Minimization
ReBeL Recursive Belief-based Learning
RL Reinforcement Learning
R-NaD Regularized Nash Dynamics
RTS Real-Time Strategy
SD-CFR Single Deep Counterfactual Regret Minimization
SPAG Self-Play of Adversarial Game
SPIN Self-Play fIne-tuNing
TSP Traveling Salesman Problem
URR UnRestricted-Restricted
XDO Extensive-Form Double Oracle